OpenStudy (natsent06):

WILL GIVE MEDAL PLEASE HELP NEED ANSWER FAST!!!! can someone help me i put the question in my calculator and change the variable each time to one of the answer choices but none of them work!! Here is the equation: Find the value of a which will make (7/8) times a = 14/16 a TRUE statement. And the answer choices: A) 2/2 B) 8/8 C) 7/7 D) 15/15

OpenStudy (natsent06):

@myininaya can you help me

OpenStudy (natsent06):

@iGreen can u help me

OpenStudy (natsent06):

(x + y) × 1/(x + y) = ?

OpenStudy (igreen):

You have to multiply 7/8 to all the options and see which one gets 14/16.

OpenStudy (natsent06):

i did but my calculator just says he answer is 7/8 every time

OpenStudy (natsent06):

(x + y) times 1/(x + y) = ?

OpenStudy (igreen):

Yeah..same here.. 7/8 = 14/16 Is this 1 choice only?

OpenStudy (natsent06):

yea

OpenStudy (igreen):

7 * 2 = 14 8 * 2 = 16 So 7/8 = 14/16

OpenStudy (natsent06):

(x + y) times 1/(x + y) = ? can u help me answer this too

OpenStudy (igreen):

\((x+y) · \dfrac{1}{(x+y)}\) Is this it?

OpenStudy (natsent06):

Select the TRUE statement. A) a divided by a = 1 B) b divided by b = 1 C) Both A and B D) Neither A nor B

OpenStudy (natsent06):

@iGreen

OpenStudy (igreen):

C.

OpenStudy (igreen):

They're both the same.